# An equivalence in random matrix and tensor models via a dually weighted intermediate field representation

### v1

@@ 5.2 The self-adjoint two-tensor model
The potentials $V$ and $Y$ place us in the genuine tensor-model setting: unlike ordinary self-adjoint matrix models, here the natural invariants are those under the product group $U(N)^{D}\subset U(N^{D})$, acting independently on each tensor slot. This symmetry imposes interactions to be of bubble type (5.11) and ensures that the Feynman diagrams retain their $D$-colored, multi-stranded structure [^10] It is important to emphasize that enlarging the symmetry group reduces the effective order of the model. If the symmetry becomes $U(N^{D/2})^{2}\subset U(N^{D})$, pairs of tensor indices are ‘locked’ together and transform as composite indices; the theory then behaves as a matrix model. If one imposes the full $U(N^{D})$ symmetry, all $D$ indices fuse into a single index, and the model degenerates into a vector model. Thus the degree of symmetry dictates the effective order of the theory: too much symmetry collapses a tensor model into a lower-order theory..
@@
$\langle\Phi\otimes\Psi\rangle[0,0]=-i\,\Sigma,\qquad\langle\Phi\otimes\Phi\rangle[0,0]=0,\qquad\langle\Psi\otimes\Psi\rangle[0,0]=0.$ (5.17)
This means that the only non–vanishing propagators are those connecting $\Phi$ and $\Psi$. Furthemore, if $V=0$ and $Y\neq 0$,
@@
with ${\cal N}=\int{\cal D}\Psi\;e^{-\frac{1}{2}{\rm Tr}\left((R\Psi)^{2}\right)}$. Thus, the two-tensor model (5.15) with a Gaussian potential in $\Psi$ effectively reduces to a Gaussian self-adjoint one–tensor model (5.22) in $\Phi$. This is directly analogous to the matrix-model case (4.12).
@@
$Y[R](\Psi)=-{\rm Tr}\ln\left({\mathbb{1}}^{\otimes D}-iR\Psi\right)$ (5.24)
@@
Let $\Phi\in{\rm H}({\mathbb{C}}_{N}^{\otimes D})$ be distributed according to $\mathcal{Z}_{\rm HT}[0,Y[R]]$. For any $\bm{\sigma}\in S_{n}^{D}$,
$\left\langle{\rm Tr}_{[\bm{\sigma}]}(\Phi)\right\rangle[0,Y[R]]=\sum_{\mu\in S_{n}}{\rm Tr}_{[\mu\bm{\sigma}]}(R)\;.$ (5.25)
@@
The similarities between Proposition 5.1 and Proposition 5.2 indicate that the Gaussian ($V=0$) sectors of the complex and self-adjoint tensor models agree at the level of trace-invariant observables. In light of the matrix-model case, this suggests an analogous equivalence between the complex tensor model $\mathcal{Z}_{\rm CT}$ with covariance $R$ and the self-adjoint tensor model $\mathcal{Z}_{\rm HT}$ with logarithmic potential (5.24). The next subsection shows that this equivalence extends to the partition functions $\mathcal{Z}_{\rm CT}$ and $\mathcal{Z}_{\rm HT}$ for any potential $V$.
@@ 5.3 Exact tensor equivalences via intermediate fields
In this section, we prove that that this agreement extends to the full ($V\neq 0$) partition functions and all observables (i.e., expectation values) depending only on the self-adjoint combination $\varphi\varphi^{\dagger}$.
5.3.1 Equivalence for scalar-$U(1)$ invariant potential $V(\varphi{\varphi^{\dagger}})$
@@
Consider the scalar-U(1)U(1)–invariant ${\mathbb{C}}_{N}^{\otimes D}$ random tensor model with partition function (5.8) and the ${\rm H}({\mathbb{C}}_{N}^{\otimes D})$ random tensor partition function $\mathcal{Z}_{\rm HT}[V,Y]$ (5.15). For $Y[R](\Psi)$ given by (5.24) and for any potential function $V$,
$\frac{\mathcal{Z}_{\rm CT}[V](R)}{\mathcal{Z}_{\rm CT}[0](R)}=\frac{\mathcal{Z}_{\rm HT}[V,Y[R]]}{\mathcal{Z}_{\rm HT}[0,Y[R]]}\;.$ (5.26)
@@
The integral over $\varphi$ is Gaussian and gives a determinant,
@@
The remaining integral is precisely ${\cal Z}_{\rm HT}[V,Y[R]]$, hence
${\cal Z}_{\rm CT}[V](R)={\cal N}_{\rm HT}^{-1}{\cal N}_{\rm CT}{\cal Z}_{\rm HT}[V,Y[R]]\;.$ (5.34)
Since $Y[R](0)=0$, one has ${\cal N}_{\rm HT}={\cal Z}_{\rm HT}[0,Y[R]]$, and the claim follows. ∎
@@
$\left\langle f(M^{\dagger}M)\right\rangle[V](R)=\frac{\mathcal{Z}_{\rm CT}(R)[V+\ln f]}{\mathcal{Z}_{\rm CT}[V](R)},\qquad\left\langle f(A)\right\rangle[V,Y[R]]=\frac{\mathcal{Z}_{\rm HT}[V+\ln f,\,Y[R]]}{\mathcal{Z}_{\rm HT}[V,Y[R]]}\;,$ (5.35)
@@
$\left\langle f(M^{\dagger}M)\right\rangle[V](R)=\left\langle f(A)\right\rangle[V,Y[R]].$ (5.36)
In particular, if $V=0$ and $f$ is a trace invariant of tensors, one recovers Propositions 5.1 and 5.2 as special cases.
The arguments above establish an equivalence between random ${\mathbb{C}}_{N}^{\otimes D}$ tensor models (5.8) and random ${\rm H}({\mathbb{C}}_{N}^{\otimes D})$ tensor models (5.15) with logarithmic potential (5.24). Moreover, under the additional partial-$U(N)$ invariance, one can further reduce the self-adjoint side to tensors of order 2(D1)2(D-1) as we now describe.
@@
Assuming a partial-$U(N)$ invariance, we consider a random tensor $\varphi\in{\mathbb{C}}_{N}^{\otimes D}$ distributed according to the partition function
@@
As well as the random tensors $\hat{\Phi},\hat{\Psi}\in{\rm H}(\mathbb{C}_{N}^{\otimes(D-1)})$ distributed according to the partition function
$).{\cal Z}_{\rm\hat{H}T}[\hat{V},\hat{Y}[R]]=\iint d\hat{\Phi}d\hat{\Psi}\;e^{-i{\rm Tr}(\hat{\Psi}\hat{\Phi})+\hat{Y}[R](\hat{\Psi})+\hat{V}(\hat{\Phi})}\,,\quad\hat{Y}[R](\hat{\Psi})=-{\rm Tr}\ln({\mathbb{1}}^{\otimes D}-R({\mathbb{1}}\otimes\hat{\Phi}$))\;. (5.38)
The following theorem relates these partition functions.
@@
$\frac{\mathcal{Z}_{\rm CT}[\hat{V}\circ{\rm Tr}_{(1)}](R)}{\mathcal{Z}_{\rm CT}[0](R)}=\frac{\mathcal{Z}_{\rm\hat{H}T}[\hat{V},\hat{Y}[R]]}{\mathcal{Z}_{\rm\hat{H}T}[0,\hat{Y}[R]]}\;.$ (5.39)
@@
The integral over $\varphi$ is Gaussian, therefore we compute
@@
The final integral is ${\cal Z}_{\rm\hat{H}T}[\hat{V},\hat{Y}[R]]$, so
${\cal Z}_{\rm CT}[\hat{V}\circ{\rm Tr}_{(1)}](R)={\cal N}_{\rm\hat{H}T}^{-1}{\cal N}_{\rm CT}{\cal Z}_{\rm\hat{H}T}[\hat{V},\hat{Y}[R]]\;.$ (5.47)
Since $\hat{Y}[R](0)=0$, one has ${\cal N}_{\rm\hat{H}T}={\cal Z}_{\rm\hat{H}T}[0,\hat{Y}[R]]$, which completes the proof. ∎
@@ 5.4 Examples
In the partial-$U(N)$ setting, if the propagator factorizes as $R=Q\otimes\hat{P}$ with $Q\in{\mathbb{C}}_{N}\otimes{\mathbb{C}}_{N}^{*}$ and $\hat{P}\in{\mathbb{C}}_{N}^{\otimes(D-1)}\otimes{\mathbb{C}}_{N}^{*\otimes(D-1)}$, the logarithmic potential simplifies to
${\cal Z}_{\rm\hat{H}T}[\hat{V},\hat{Y}[R]]=\iint{\cal D}\hat{\Phi}{\cal D}\hat{\Psi}\;e^{-i{\rm Tr}(\hat{\Psi}\hat{\Phi})+\sum_{k=1}^{\infty}\frac{i^{k}}{k}{\rm Tr}\left(Q^{k}\right){\rm Tr}\left((\hat{P}\hat{\Psi})^{k}\right)+\hat{V}(\hat{\Phi})}\;$, (5.49)
@@
with $\varphi\in{\mathbb{C}}_{N}^{\otimes D}$.
Moreover, if $Q=C_{2}$, one obtains the Gaussian reduction
${\cal Z}_{\rm\hat{H}T}[\hat{V},\hat{Y}[R]]=\iint{\cal D}\hat{\Phi}{\cal D}\hat{\Psi}\;e^{-i{\rm Tr}(\hat{\Psi}\hat{\Phi})-N\frac{1}{2}{\rm Tr}\left((\hat{P}\hat{\Psi})^{2}\right)+\hat{V}(\hat{\Phi})}\;$, (5.51)
@@
${\cal Z}_{\rm\hat{H}T}[\hat{V},\hat{Y}[R]]={\cal N}\int{\cal D}\hat{\Phi}\;e^{-N\frac{1}{2}{\rm Tr}\left((\hat{P}^{-1}\hat{\Phi})^{2}\right)+\hat{V}(\hat{\Phi})}\;$, (5.52)
@@
From an order–$D$ quartic pillow model to an order–$2D$ quadratic self-adjoint model.
@@
Figure 10: The three types of pillow vertices present in the self-adjoint tensor model with potential (5.55). These are labeled consistently with the three pillows in Figure 9.
Next, we consider a random complex tensor $\varphi\in\mathbb{C}_{N}^{\otimes D}$ with partition function
@@
${\cal Z}_{\rm HT}[{V}_{\rm 4},Y[R]]=\iint{\cal D}\Phi{\cal D}\Psi\;e^{-i{\rm Tr}\left(\Psi\Phi\right)-{\rm Tr}\ln\left({\mathbb{1}}^{\otimes D}-iR\Psi\right)+V_{4}(\Phi)}\;.$ (5.57)
@@
$\frac{{\cal Z}_{\rm CT}[V_{4}](R)}{{\cal Z}_{\rm CT}[0](R)}=\frac{{\cal Z}_{\rm HT}[V_{4},Y[R]]}{{\cal Z}_{\rm HT}[0,Y[R]]}\;.$ (5.58)
In this example the complex quartic pillow interaction on the order–$D$ tensor is mapped to a purely quadratic interaction on the order–$2D$ self–adjoint tensor, providing a particularly simple illustration of how bubble structure changes under the equivalence.
From an order–$D$ quartic pillow model to an order–2(D1)2(D-1) quadratic self-adjoint model.
@@
$V_{\rm 4}(\varphi{\varphi^{\dagger}})=N\frac{\lambda}{2}\sum_{\bm{a},\bm{b}}\left(\varphi^{{\bm{b}}_{\hat{1}}{\bm{a}}}{\varphi^{\dagger}}_{\bm{a}}\varphi^{{\bm{a}}_{\hat{1}}\bm{b}}{\varphi^{\dagger}}_{\bm{b}}+\sum_{c=2}^{D}\varphi^{{\bm{a}}_{\hat{c}}\bm{b}}{\varphi^{\dagger}}_{\bm{a}}\varphi^{{\bm{b}}_{\hat{c}}{\bm{a}}}{\varphi^{\dagger}}_{\bm{b}}\right)\;,$ (5.59)
@@
Next, we consider a random complex tensor $\varphi\in\mathbb{C}_{N}^{\otimes D}$ with partition function
@@
${\cal Z}_{\rm\hat{H}T}[\hat{V}_{\rm 4},Y[R]]=\iint_{{\rm H}({{\mathbb{C}}_{N}^{\otimes(D-1)}})^{2}}{\cal D}\hat{\Phi}{\cal D}\hat{\Psi}\;e^{-i{\rm Tr}(\hat{\Psi}\hat{\Phi})-{\rm Tr}\ln\left({\mathbb{1}}^{\otimes D}-iR({\mathbb{1}}\otimes\hat{\Psi})\right)+\hat{V}_{4}(\Phi)}\;.$ (5.62)
@@
$\frac{{\cal Z}_{\rm CT}[V_{4}](R)}{{\cal Z}_{\rm CT}[0](R)}=\frac{{\cal Z}_{\rm\hat{H}T}[\hat{V}_{4},Y[R]]}{{\cal Z}_{\rm\hat{H}T}[0,Y[R]]}\;.$ (5.63)
In this example the complex quartic pillow interaction on the order–$D$ tensor is mapped to a purely quadratic interaction on the order–2(D1)2(D-1) self–adjoint tensor, providing a particularly simple illustration of how bubble structure changes under the equivalence.
In particular, if $R=C_{2}\otimes\hat{P}$, the self-adjoint two-tensor model becomes a one-tensor model with partition function
${\cal Z}_{\rm\hat{H}T}[\hat{V}_{4},\hat{Y}[C_{2}\otimes\hat{P}]]=\iint_{{\rm H}({{\mathbb{C}}_{N}^{\otimes(D-1)}})}{\cal D}\hat{\Phi}\;e^{-\frac{N}{2}{\rm Tr}((\hat{P}^{-1}\hat{\Psi})^{2})+\hat{V}_{4}(\hat{\Psi})}\;$, (5.64)

### v2

@@ 6.2 The self-adjoint two-tensor model
The potentials $V$ and $Y$ place us in the genuine tensor-model setting: unlike ordinary self-adjoint matrix models, here the natural invariants are those under the product group $U(N)^{D}\subset U(N^{D})$, acting independently on each tensor factor $\mathbb{C}_{N}$. This symmetry imposes interactions to be of bubble type (6.11) and ensures that the Feynman diagrams retain their $D$-colored, multi-stranded structure [^12] It is important to emphasize that enlarging the symmetry group reduces the effective order of the model. If the symmetry becomes $U(N^{D/2})^{2}\subset U(N^{D})$, pairs of tensor indices are ‘locked’ together and transform as composite indices; the theory then behaves as a matrix model. If one imposes the full $U(N^{D})$ symmetry, all $D$ indices fuse into a single index, and the model degenerates into a vector model. Thus the degree of symmetry dictates the effective order of the theory: too much symmetry collapses a tensor model into a lower-order theory..
@@
$\langle\Phi\otimes\Psi\rangle[0,0]=-i\,\Sigma,\qquad\langle\Phi\otimes\Phi\rangle[0,0]=0,\qquad\langle\Psi\otimes\Psi\rangle[0,0]=0\,$, (6.17)
where $\Sigma^{\bm{a}\bm{c}}_{\bm{b}\bm{d}}=\delta^{\bm{a}}_{\bm{d}}\delta^{\bm{c}}_{\bm{b}}$ as defined in Section 2. This means that the only non–vanishing propagators are those connecting $\Phi$ and $\Psi$. This propagator is illustrated in Figure 14.
Figure 14: Stranded graph representation of the propagator of the theory defined by the partition function (6.15) as given in (6.17).
Furthemore, if $V=0$ and $Y\neq 0$,
@@
with ${\cal N}=\int{\cal D}\Psi\;e^{-\frac{1}{2}{\rm Tr}\left((R\Psi)^{2}\right)}$. Thus, the two-tensor model (6.15) with a Gaussian potential in $\Psi$ and an arbitrary $V(\Phi)$ effectively reduces to a self-adjoint one–tensor model (6.22) in $\Phi$ with $V(\Phi)$. This is directly analogous to the matrix-model case (5.12).
@@
$Y_{\ln}[R](\Psi)=-{\rm Tr}\ln\left({\mathbb{1}}^{\otimes D}-iR\Psi\right)=\sum_{k=1}^{\infty}\frac{i^{k}}{k}{\rm Tr}((R\Psi)^{k})$ (6.24)
@@
Let $\Phi\in{\rm H}({\mathbb{C}}_{N}^{\otimes D})$ be distributed according to $\mathcal{Z}_{\rm HT}[0,Y_{\ln}[R]]$. For any $\bm{\sigma}\in S_{n}^{D}$,
$\left\langle{\rm Tr}_{[\bm{\sigma}]}(\Phi)\right\rangle[0,Y_{\ln}[R]]=\sum_{\mu\in S_{n}}{\rm Tr}_{[\mu\bm{\sigma}]}(R)\;.$ (6.25)
@@
The similarities between Proposition 6.1 and Proposition 6.2 indicate that the Gaussian ($W=0$ for Proposition 6.1 and $V=0$ for Proposition 6.2) sectors of the complex and self-adjoint tensor models agree at the level of trace-invariant observables. In light of the matrix-model case, this suggests an analogous equivalence between the complex tensor model $\mathcal{Z}_{\rm CT}$ with covariance $R$ and the self-adjoint tensor model $\mathcal{Z}_{\rm HT}$ with logarithmic potential (6.24). The next subsection shows that this equivalence extends to the partition functions $\mathcal{Z}_{\rm CT}$ and $\mathcal{Z}_{\rm HT}$ for any potential $V$.
@@ 6.3 Exact tensor equivalences via intermediate fields
In this section, we show that this agreement extends to the full ($V\neq 0$) partition functions and to all observables (i.e. expectation values) depending only on the self-adjoint combination $\varphi\varphi^{\dagger}$. Similarly to the matrix model equivalence, this is done through an intermediate field representation. In this formulation, the self-adjoint tensor $\Phi$ appears as an intermediate effective field, while a second self-adjoint tensor $\Psi$ appears as a dually-weighted intermediate field in which the logarithmic interaction, deformed by the covariance inherited from the propagator of the complex model, gives rise to the characteristic dually weighted structure.
6.3.1 Equivalence for a scalar-$U(1)$ invariant potential $V(\varphi{\varphi^{\dagger}})$
Assuming a scalar-$U(1)$ invariance, we consider a complex tensor $\varphi\in{\mathbb{C}}_{N}^{\otimes D}$ distributed according to the partition function
${\cal Z}_{\rm CT}[V](R)=\iint{\cal D}\varphi^{\dagger}{\cal D}\varphi\;e^{-\varphi^{\dagger}R^{-1}\varphi+V(\varphi\varphi^{\dagger})}\;,$ (6.26)
with the complex field $\varphi$ carrying the potential $V(\varphi\varphi^{\dagger})$. We also consider self-adjoint tensors $\Phi,\Psi\in{\rm H}(\mathbb{C}_{N}^{\otimes D})$ distributed according to
${\cal Z}_{\rm HT}[V,Y_{\ln}[R]]=\iint{\cal D}\Phi{\cal D}\Psi\;e^{-i{\rm Tr}\left(\Phi\Psi\right)+Y_{\ln}[R](\Psi)+V(\Phi)}\;,\quad{Y}_{\ln}[R](\Psi)=-{\rm Tr}\ln({\mathbb{1}}^{\otimes D}-iR\Psi)\;.$ (6.27)
In this formulation, $\Phi$ plays the role of an intermediate effective field, carrying an effective potential $V(\Phi)$, while $\Psi$ appears as a dually-weighted intermediate field, carrying a dually-weighted logarithmic potential $Y_{\ln}[R](\Psi)$. The following theorem establishes the equivalence between these partition functions.
@@ 6.3.1 Equivalence for a scalar-$U(1)$ invariant potential $V(\varphi{\varphi^{\dagger}})$
Consider the scalar-U(1)U(1)–invariant ${\mathbb{C}}_{N}^{\otimes D}$ random tensor model with partition function (6.26) and the ${\rm H}({\mathbb{C}}_{N}^{\otimes D})$ random tensor partition function $\mathcal{Z}_{\rm HT}[V,Y_{\ln}[R]]$ (6.27). For any potential function $V$,
$\frac{\mathcal{Z}_{\rm CT}[V](R)}{\mathcal{Z}_{\rm CT}[0](R)}=\frac{\mathcal{Z}_{\rm HT}[V,Y_{\ln}[R]]}{\mathcal{Z}_{\rm HT}[0,Y_{\ln}[R]]}\;$ (6.28)
as formal power series.
@@
For sufficiently well-behaved interactions $V$, one may exchange the order of integration and perform the integral over $\varphi$ first. More generally, for any $V$, this operation can always be performed at the level of formal power series. The integral over $\varphi$ is Gaussian, and can therefore be computed explicitly,
@@
The remaining integral is precisely ${\cal Z}_{\rm HT}[V,Y_{\ln}[R]]$, hence
${\cal Z}_{\rm CT}[V](R)={\cal N}_{\rm HT}^{-1}{\cal N}_{\rm CT}{\cal Z}_{\rm HT}[V,Y_{\ln}[R]]\;.$ (6.36)
Since $Y_{\ln}[R](0)=0$, one has ${\cal N}_{\rm HT}={\cal Z}_{\rm HT}[0,Y_{\ln}[R]]$, and the claim follows. ∎
Remark. As in the matrix case, the equivalence is carried at the level of formal power series, since for general choices of interaction the tensor integrals may fail to be absolutely convergent or may only be defined in a distributional sense. For suitable potentials, however, the equivalence extends to well-defined tensor integrals.
One explicit example is given by mapping the sextic complete interaction model to a cubic interaction model as shown in Figure 16, which includes the vertices shown in Figure 15. The sextic model is explored further in (6.73).
Figure 15: (a) A sextic interaction vertex that can be present in the potential $W$ of the partition function (6.1). (b) The $i\varphi^{\dagger}\Psi\varphi$ vertex in (6.33). (c) An example of a cubic potential of the intermediate effective field $\Phi$ which may be present in $V$ in (4.6). (d) The cubic potential of the dually-weighted intermediate field $\Psi$, which appears as the third order term in the expansion of the potential ${\rm Tr}\ln\left({\mathbb{1}}^{\otimes D}-iR\Psi\right)$ in (4.6).
Figure 16: An example of graphs whose amplitudes are equal in the different tensor models. Those graphs are composed by the vertices shown in Figure 15 and the propagators shown in Figure 12 and Figure 14. (a) A graph that can be generated by the partition function (6.1); (b-c) Two representations of the same graph that can be generated by the partition function (6.33); (d) A graph that can be generated by the partition function (4.6).
@@
$\left\langle f(\varphi\varphi^{\dagger})\right\rangle[V](R)=\frac{\mathcal{Z}_{\rm CT}(R)[V+\ln f]}{\mathcal{Z}_{\rm CT}[V](R)},\qquad\left\langle f(\Psi)\right\rangle[V,Y[R]]=\frac{\mathcal{Z}_{\rm HT}[V+\ln f,\,Y_{\ln}[R]]}{\mathcal{Z}_{\rm HT}[V,Y_{\ln}[R]]}\;,$ (6.37)
@@
$\left\langle f(\varphi\varphi^{\dagger})\right\rangle[V](R)=\left\langle f(\Phi)\right\rangle[V,Y_{\ln}[R]].$ (6.38)
In particular, if $V=0$ and $f$ is a trace invariant of tensors, one recovers Propositions 6.1 and 6.2 as special cases.
The arguments above establish an equivalence between random ${\mathbb{C}}_{N}^{\otimes D}$ tensor models (6.8) and random ${\rm H}({\mathbb{C}}_{N}^{\otimes D})$ tensor models (6.15) with logaritmic potential (6.24). Moreover, under the additional partial-$U(N)$ invariance, one can further reduce the self-adjoint side to tensors of order 2(D1)2(D-1) as we now describe.
@@
Assuming partial-$U(N)$ invariance, we consider a complex tensor $\varphi\in{\mathbb{C}}_{N}^{\otimes D}$ distributed according to the partition function
@@
We also consider self-adjoint tensors $\hat{\Phi},\hat{\Psi}\in{\rm H}(\mathbb{C}_{N}^{\otimes(D-1)})$ distributed according to
$).{\cal Z}_{\rm\hat{H}T}[\hat{V},\hat{Y}_{\ln}[R]]=\iint d\hat{\Phi}d\hat{\Psi}\;e^{-i{\rm Tr}(\hat{\Psi}\hat{\Phi})+\hat{Y}_{\ln}[R](\hat{\Psi})+\hat{V}(\hat{\Phi})}\,,\quad\hat{Y}_{\ln}[R](\hat{\Psi})=-{\rm Tr}\ln({\mathbb{1}}^{\otimes D}-iR({\mathbb{1}}\otimes\hat{\Psi}$))\;. (6.40)
In this formulation, $\hat{\Phi}$ plays the role of an intermediate effective field, while $\hat{\Psi}$ appears as a dually-weighted intermediate field. The following theorem establishes the equivalence between these partition functions.
@@
$\frac{\mathcal{Z}_{\rm CT}[\hat{V}\circ{\rm Tr}_{(1)}](R)}{\mathcal{Z}_{\rm CT}[0](R)}=\frac{\mathcal{Z}_{\rm\hat{H}T}[\hat{V},\hat{Y}_{\ln}[R]]}{\mathcal{Z}_{\rm\hat{H}T}[0,\hat{Y}_{\ln}[R]]}\;$ (6.41)
as formal power series.
@@
For sufficiently well-behaved interactions, one may exchange the order of integration and perform the integral over $\varphi$ first. More generally, this step can be understood at the level of formal power series. The integral over $\varphi$ is Gaussian, and can therefore be computed explicitly,
@@
The final integral is ${\cal Z}_{\rm\hat{H}T}[\hat{V},\hat{Y}_{\ln}[R]]$, so
${\cal Z}_{\rm CT}[\hat{V}\circ{\rm Tr}_{(1)}](R)={\cal N}_{\rm\hat{H}T}^{-1}{\cal N}_{\rm CT}{\cal Z}_{\rm\hat{H}T}[\hat{V},\hat{Y}_{\ln}[R]]\;.$ (6.49)
Since $\hat{Y}_{\ln}[R](0)=0$, one has ${\cal N}_{\rm\hat{H}T}={\cal Z}_{\rm\hat{H}T}[0,\hat{Y}_{\ln}[R]]$, which completes the proof. ∎
Remark. The same considerations as in the matrix case and in Theorem 6.3 apply: the equivalence is understood at the level of formal power series in general, and extends to convergent tensor integrals whenever the latter are well-defined.
@@ 6.4 Examples
In the partial-$U(N)$ setting, if the propagator factorizes as $R=Q\otimes\hat{P}$ with $Q\in{\mathbb{C}}_{N}\otimes{\mathbb{C}}_{N}^{*}$ and $\hat{P}\in{\mathbb{C}}_{N}^{\otimes(D-1)}\otimes{\mathbb{C}}_{N}^{*\otimes(D-1)}$, the logarithmic potential (6.40) simplifies to
${\cal Z}_{\rm\hat{H}T}[\hat{V},\hat{Y}_{\ln}[R]]=\iint{\cal D}\hat{\Phi}{\cal D}\hat{\Psi}\;e^{-i{\rm Tr}(\hat{\Psi}\hat{\Phi})+\sum_{k=1}^{\infty}\frac{i^{k}}{k}{\rm Tr}\left(Q^{k}\right){\rm Tr}\left((\hat{P}\hat{\Psi})^{k}\right)+\hat{V}(\hat{\Phi})}\;$, (6.51)
@@
with $\varphi\in{\mathbb{C}}_{N}^{\otimes D}$. Moreover, if $Q=C_{2}$ (see (3.6)), one obtains the Gaussian reduction
${\cal Z}_{\rm\hat{H}T}[\hat{V},\hat{Y}_{\ln}[R]]=\iint{\cal D}\hat{\Phi}{\cal D}\hat{\Psi}\;e^{-i{\rm Tr}(\hat{\Psi}\hat{\Phi})-N\frac{1}{2}{\rm Tr}\left((\hat{P}\hat{\Psi})^{2}\right)+\hat{V}(\hat{\Phi})}\;$, (6.53)
@@
${\cal Z}_{\rm\hat{H}T}[\hat{V},\hat{Y}_{\ln}[R]]={\cal N}\int{\cal D}\hat{\Phi}\;e^{-N\frac{1}{2}{\rm Tr}\left((\hat{P}^{-1}\hat{\Phi})^{2}\right)+\hat{V}(\hat{\Phi})}\;$, (6.54)
@@
Equivalences for trace-invariant potentials.
[…]
More specific examples of these trace-invariant potentials can be found in
  • (i)
  • (ii)
See also the corresponding Figures for (i) Figures 17 and 18, and (ii) Figures 19 and 20 for the graphic illustrations of these trace-invariant potentials.
(i) From an order–$D$ quartic pillow model to an order–$2D$ quadratic self-adjoint model.
@@
Figure 18: The three types of quadratic vertices of the intermediate effective field $\Phi$ present in the self-adjoint tensor model with potential (6.62). These are labelled consistently with the three quartic pillows in Figure 17.
We consider a random complex tensor $\varphi\in\mathbb{C}_{N}^{\otimes D}$ with partition function
@@
${\cal Z}_{\rm HT}[{V}_{\rm 4},{Y}_{\ln}[R]]=\iint{\cal D}\Phi{\cal D}\Psi\;e^{-i{\rm Tr}\left(\Psi\Phi\right)-{\rm Tr}\ln\left({\mathbb{1}}^{\otimes D}-iR\Psi\right)+V_{4}(\Phi)}\;.$ (6.64)
@@
$\frac{{\cal Z}_{\rm CT}[V_{4}](R)}{{\cal Z}_{\rm CT}[0](R)}=\frac{{\cal Z}_{\rm HT}[V_{4},{Y}_{\ln}[R]]}{{\cal Z}_{\rm HT}[0,{Y}_{\ln}[R]]}\;.$ (6.65)
In this example the complex quartic pillow interaction on the order–$D$ tensor is mapped to a purely quadratic interaction on the order–$2D$ self–adjoint tensor.
(ii) From an order–$D$ quartic pillow model to an order–2(D1)2(D-1) quadratic self-adjoint model.
@@
$V_{\rm 4}(\varphi{\varphi^{\dagger}})=N\frac{\lambda}{2}\sum_{\bm{a},\bm{b}}\left({\varphi^{\dagger}}_{\bm{a}}\varphi^{{\bm{b}}_{\hat{1}}{\bm{a}}}{\varphi^{\dagger}}_{\bm{b}}\varphi^{{\bm{a}}_{\hat{1}}\bm{b}}+\sum_{c=2}^{D}{\varphi^{\dagger}}_{\bm{a}}\varphi^{{\bm{a}}_{\hat{c}}\bm{b}}{\varphi^{\dagger}}_{\bm{b}}\varphi^{{\bm{b}}_{\hat{c}}{\bm{a}}}\right)\;,$ (6.66)
@@
We consider a random complex tensor $\varphi\in\mathbb{C}_{N}^{\otimes D}$ with partition function
@@
${\cal Z}_{\rm\hat{H}T}[\hat{V}_{\rm 4},{Y}_{\ln}[R]]=\iint_{{\rm H}({{\mathbb{C}}_{N}^{\otimes(D-1)}})^{2}}{\cal D}\hat{\Phi}{\cal D}\hat{\Psi}\;e^{-i{\rm Tr}(\hat{\Psi}\hat{\Phi})-{\rm Tr}\ln\left({\mathbb{1}}^{\otimes D}-iR({\mathbb{1}}\otimes\hat{\Psi})\right)+\hat{V}_{4}(\Phi)}\;.$ (6.69)
@@
$\frac{{\cal Z}_{\rm CT}[V_{4}](R)}{{\cal Z}_{\rm CT}[0](R)}=\frac{{\cal Z}_{\rm\hat{H}T}[\hat{V}_{4},{Y}_{\ln}[R]]}{{\cal Z}_{\rm\hat{H}T}[0,\ {Y}_{\ln}[R]]}\;.$ (6.70)
In this example the complex quartic pillow interaction on the order–$D$ tensor is mapped to a purely quadratic interaction on the order–2(D1)2(D-1) self–adjoint tensor.
In particular, if $R=C_{2}\otimes\hat{P}$, the self-adjoint two-tensor model (6.69) becomes a one-tensor model with partition function as seen in (6.54),
${\cal Z}_{\rm\hat{H}T}[\hat{V}_{4},\hat{Y}_{\ln}[C_{2}\otimes\hat{P}]]=\iint_{{\rm H}({{\mathbb{C}}_{N}^{\otimes(D-1)}})}{\cal D}\hat{\Phi}\;e^{-\frac{N}{2}{\rm Tr}((\hat{P}^{-1}\hat{\Phi})^{2})+\hat{V}_{4}(\hat{\Phi})}\;$, (6.71)